Tags:
create new tag
view all tags

Shift 20120524 LC LN AA

Shift plan

  • Load LW optics.
  • OTR data - polariser data, during polariser analysis do LW spatial alignment.
  • Quick OTR quad scan to find minimum σy. Also get σx for this beam.
  • Try various skew quads to see which one rotates the beam significantly on the OTR screen.
  • Tune for beam steering and background on LW detector - ideally < 7000 counts.
  • Timing.
  • Collisions search. Preliminary vertical/horizontal scans to find maximum of signal.
  • Take long background and peak logs.
  • Laser energy/signal correlation.
  • Dispersion measurement - find out how to display the calculated dispersion at our location - and correction.
  • Skew scan if good magnet found previously.
  • Detailed long vertical and horizontal scans for full data collection.
  • Vertical slices at various points through horizontal aspect of smallest beam.
  • Quad scans - QM14FF first, then maybe another quad?
  • Matching OTR quad scans.

PreShift Log

  • Updated pedestal for ADCs with detector switched on, but no voltage and in addition laser closed for laserpd.

  • 1000 pulse log after ADC pedestal update and with detector and laser off. 20120523_1955_log.dat

Shift

02:00

Tuning for background - odd orbit, hard to get background under 10,000 with 10dB of attenuation.

03:25 OTR scans

Rotation angle 4.32 degrees. Polariser scan min at 19.38 degrees. Set to 19.38. Try more OTR quad scans but beam charge very low so beam barely visible.

05:30

Laser running very low energy, bad signal to noise. Phase scan 20120524_0546_lws.dat, centre 3.05V. Background ~ 10,000 counts with 3dB attenuation.

06:00

Warning septum leak, beam off to investigate.

06:15

Not water leak, sensor is broken, beam back on. Vertical scan 20120524_0615_lws.dat, cetnre 809. Horizontal scan 20120524_0620_lws.dat, centre at 490. Move to this position.

06:50

Signal very low. Try different telescope settings - 5mm (large beam axis diameter 14mm, small beam axis 10mm) and 16mm (large beam axis diameter 10.5mm, small beam axis diameter 7.5mm). All telescope measurements from 17_5_2012.

5mm scan 20120524_0648_lws.dat 16mm scan 20120524_0652_lws.dat

No signal, return to 12.7mm (original telescope setting).

08:00

Signal decreasing, laser energy on post IP meter right down to 30 - 45mJ. Check laser, regen ~ 9mJ but not significantly worse than start of shift. Take access to check experiment - input window badly damaged, output window slightly damaged, lens fine. Gate valves around LWIP closed.

09:00

Discussion with Terunuma-san about operating accelerator, replacing window or moving chamber. Decided to operate machine and monitor vacuum closely (no sign of leak yet, window 56mm diameter, 12.7mm thick).

09.20

Beam on.

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r6 - 24 May 2012 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback